MOM’s report from G.CECCHET (abridged). MOM’s report First half of November we had normal operations with some data taking for Marco, Yordan, Maurizio.
Published byModified over 5 years ago
Presentation on theme: "MOM’s report from G.CECCHET (abridged). MOM’s report First half of November we had normal operations with some data taking for Marco, Yordan, Maurizio."— Presentation transcript:
MOM’s report First half of November we had normal operations with some data taking for Marco, Yordan, Maurizio. DAQ had some instabilities temporarily patched waiting for JSG to reshuffle the electronics (just doing so). Target: some problems at the beginning due to an incorrect timing (solved)
MOM’s report Decay solenoid: working normally until the small air compressor failed, causing the shutdown of the Linde compressor and the warming up of DS. The air compressor has been repaired and the cool down restarted. Full operation should be achieved soon.
MOM’s report Conventional magnets: normal behaviour except for a temporarily (and unexplained) Q6 failure. Normal operation have been resumed after one day. PID detectors: after changing a pm with Maurizio in TOF0, both TOF’s behaved normally, except for a problem to the HV system.
MOM’s report HV system: the SY527 had a failure in one of the cards controlling the HV for TOF1. The spare module arrived with TOF2 and the new stand for TOF2/KL, plus a second spare. TOF2 installation: on Nov 23 rd we started the installation of the new stand for KL/TOF2. Few pics follow:
MOM’s report The new stand is a movable one and the relative position of TOF2/KL can be changed to a certain extent. All the pm’s of TOF2 have been tested, the HV cables laid and connected to SY527 and signal cables brought for the detector to the MLCR, ready for JSG to be integrated in the readout electronics.
MOM’s report Minicom: all the HV values have been restored in the Minicom program. The detectors’ HV have been grouped into logical names and the default values are now in paper form in the folder in MLCR and there are pointers to the Pavia’s database (see next News).
MOM’s report Administration: an agreement has been reached with ISIS people for beam time requests. By now everyone should be aware of the procedure and that will be useful to have some coordination in our work.
MOM’s report Caveats: I have few comments on some security issues. The daily check of the water level on the roof, when raining, could be dangerous due to the slippery surface of the metal stair and the wooden path. Is it possible to have an electronic measure of that level and the readout of the conductivity meter?
MOM’s report Beam blocker: everybody agree that it is not failsafe. I understand that it is now difficult (and expensive) to change it. But is it possible to redesign the blocking system in order to make it accessible to everybody?
MOM’s report And now I want to thank everybody for the support I received during this period. Particularly Marco who has been a patient tutor to me. In the name of the TOF/KL collaborators, I want to thank Willie and his team for all the work they have done for the installation of TOF2. And now……